OpenStudy (anonymous):

Is this is an empty solution?? √2t+4=-4 Thx!!!

OpenStudy (experimentx):

if t is inside sqrt then, t = 32

OpenStudy (anonymous):

yes, 2t+4 are inside the sqrt.

OpenStudy (experimentx):

then your solution would be 6

OpenStudy (anonymous):

\[\sqrt{2t+4}=-4\]

OpenStudy (anonymous):

no solution, of course in \[\mathbb{R}\]

OpenStudy (anonymous):

Thanks!! Cinar..

OpenStudy (anonymous):

\[\sqrt{x}\ge0\]

OpenStudy (anonymous):

np